OpenStudy (anonymous):

what is the multiplicative inverse of -1/9

myininaya (myininaya):

Let b be the multiplicative inverse of a such ab=1

myininaya (myininaya):

S the multiplicative inverse of -1/9 is b and a is -1/9 Solve for the multiplicative inverse of -1/9

myininaya (myininaya):

In other words, solve for b

myininaya (myininaya):

ab=1 a*b=1 -1/9 * b=1 <---solve for b Can you do this?

OpenStudy (anonymous):

you lost me big time e.0???

myininaya (myininaya):

Your question ask us to find the multiplicative inverse of a I'm calling that multiplicative inverse b. The multiplicative identity is 1. If a and b are multiplicative inverses of each other, then the product of a and b is 1.

myininaya (myininaya):

ab=1 -1/9 * b=1 You need to solve this for b in other words.

myininaya (myininaya):

If you don't know how to solve equations, then you can always remember to take the opposite of the reciprocal. Just so you know 0 doesn't have a multiplicative inverse.

myininaya (myininaya):

What is the reciprocal of -1/9?

myininaya (myininaya):

To find the reciprocal just switch the numerator and denominator for each other.

myininaya (myininaya):

Do you know what I mean?

Parth (parthkohli):

Opposite of the reciprocal? Multiplicative inverse IS reciprocal.

myininaya (myininaya):

Oops. I'm thinking of lines that are perpendicular. I'm sorry.

myininaya (myininaya):

Just flip it. lol.

Parth (parthkohli):

:)

myininaya (myininaya):

But still ab=1 then a and b are multiplicative inverses of each other